140625 LML Observations

Isaac arrived at the LML site at 10:30. The doors to the fog collector had opened although no visible sample had been obtained. The top plate, drip tube assembly, and the strand frames were removed for cleaning. The body and fog collector components were then sprayed initially with 409 followed by thorough rinsing with DI water. The system was then reassembled for collection during the following day.

Leave a Reply

Your email address will not be published. Required fields are marked *